The perimeter of a rectangle is 60 meters. The length of the rectangle is 6 meters less than twice the width. Find the length and with of the rectangle.
Answer by ewatrrr(24785)


Hi,
Let x and (2x-6) represent the width and length respectively
P = 2*l + 2*w ***
60m = 2(2x-6) + 2x
Solving for x
60m = 4x - 12m + 2x
72m = 6x
x = 12m, the width. The length is 18m. (2*12 - 6)
CHECKING our Answer***
2*12m + 2*18m = 24m + 36m = 60m